OpenStudy (anonymous):

Solve for x. 1/3(9-6x)+x=7+2x-2

OpenStudy (aaronq):

move all the numbers to one side and all the x's to the other

OpenStudy (anonymous):

I'm really bad with fractions

OpenStudy (aaronq):

1/3(9-6x)+x=7+2x-2 9-6x-2x=5*3 -6x-2x=15-9

OpenStudy (aaronq):

there's no fractions really, it's (9-6x) divided by 3, when you bring it over you get rid of the brackets. can you finish it?

OpenStudy (anonymous):

do I combine the fractions?

OpenStudy (anonymous):

the negative ones?

OpenStudy (anonymous):

you can always multiply both sides by 3 to get rid of the fraction

OpenStudy (aaronq):

are you referring to the the x's (unknown/variables) as fractions?

OpenStudy (anonymous):

the 1/3

OpenStudy (anonymous):

1/3(9-6x)+x=7+2x-2 multiply both sides by 3 to get rid of 1/3 then :P 9-6x=3*(7+2x-2)

OpenStudy (aaronq):

yeah, just move it over to the other side by performing the inverse function which would be multiplication

OpenStudy (anonymous):

Will it be -2/3..?

OpenStudy (johnweldon1993):

I believe you missed an 'x' @timo86m 1/3(9 - 6x) + x = 7 + 2x - 2 subtract x from both sides first 1/3(9 - 6x) = 7 + x - 2 Now multiply both sides by 3 9-6x = 3(7 + x - 2) 9 - 6x = 21 + 3x - 6 9 - 6x = 15 + 3x 9 = 15 + 9x 9x = -6 x = -6/9 x = -2/3 Yes good job @cynthia_franco

OpenStudy (anonymous):

D'awh, thank you!

OpenStudy (anonymous):

sorry about that